Suppose a bank has $100,000 in checking account deposits with no excess reserves and the required reserve ratio is 10 percent. If the Federal Reserve raises the required reserve ratio to 12 percent, then the bank will now have excess reserves of

A) $12,000.

B) $0.

C) -$2,000.

D) -$12,000.
Answer #1

Answer

required reserves =deposits * reserve ratio

before the change in the reserve ratio

required reserves =100000*0.1=$10000

after the change

required reserves =100000*0.12=$12000

but the available reserves is $10000 as there are no excess reserves

so the

excess reserves =10000-12000=-$2000

so the excess reserves are -$2000

option D